# Q-ball polarization - a smooth path to oscillons

### v1

@@ 1 Motivation
Oscillons and Q-balls are examples of non-topological, solitonic-like excitations. At first glance, they are rather different objects. The Q balls are stationary solutions that periodically oscillate in the target complex space [1, 2]. They are stable due to the $U(1)$ Noether charge. Oscillons, on the contrary, are quasi-periodic solutions for which the energy density oscillates in time [3, 4, 5]. The fundamental frequency of the oscillations is, however, accompanied by a tower of frequencies. Oscillons exist even in real field theories and do not carry any charge. Thus, their existence is not related, in an obvious manner, to any conservation law. In fact, they are not ultimately stable, but eventually decay, sometimes extremely slowly [6].
@@
where $\omega<m$ is the internal frequency. Typically, $\omega<m$, where $m$ is the mass threshold. Often, there is also a lower bound for $\omega$. Next, purely real, small amplitude fundamental (i.e., without modulations) oscillons are approximated by
@@ 5 Conclusions
We showed that the oscillons and Q-balls in the $\phi^{6}$ model are approximately ”dual” objects. One can be approximately expressed by the other. Namely, a Q-ball can be viewed as a superposition of two oscillons, or equivalently, the oscillon can be treated as a Q-ball and anti-Q-ball pair. In a sense, both represent different types of ”polarization” of the complex field.

### v2

@@ 1 Motivation
Oscillons and Q-balls are examples of non-topological, solitonic-like excitations. At first glance, they are rather different objects. The Q balls are stationary solutions that periodically oscillate in the target complex space [1, 2]. They are stable due to the $U(1)$ Noether charge. Oscillons, on the contrary, are quasi-periodic solutions for which the energy density oscillates in time [3, 4, 5]. The fundamental frequency of the oscillations is, however, accompanied by a tower of frequencies. Oscillons exist even in real field theories and do not carry any charge. Thus, their existence is not related, in an obvious manner, to any conservation law. In fact, they are not ultimately stable, but eventually decay, sometimes extremely slowly [6, 7, 8, 9].
@@
with $\omega<m$, where $\omega$ is the internal frequency and $m$ is the mass threshold. Often, there is also a lower bound for $\omega$. Next, purely real, small amplitude fundamental (i.e., without modulations) oscillons are approximated by
@@ 5 Conclusions
We showed that the oscillons and Q-balls in the complex $\phi^{6}$ model are approximately ”dual” objects. One can be approximately expressed by the other. Namely, a Q-ball can be viewed as a superposition of two oscillons, or equivalently, the oscillon can be treated as a Q-ball and anti-Q-ball pair. In a sense, both represent different types of ”polarization” of the complex field.
@@
$Q$-balls as well as oscillons appear in various cosmological scenarios. They can be produced in the early universe, especially at the end of inflation [23, 24, 25, 26]. They may contribute to the dark matter sector [27, 28]. They can also play a role in baryogenesis [29, 30]. It would be very interesting to understand the physical consequence of the fact that $Q$ balls and oscillons seem to be smoothly deformable one into the other.